what is v in the following equation 14v+6=2(5+7v)-4
Answer by nyc_function(2741)

14v+6 =2(5+7v)-4

14v + 6 = 10 + 14v - 4

I can see right away this linear equation has no solution because 14v will

cancel out with -14v.

14v + 6 = 14v + 6

The only thing that can be said about this equation as a whole is that one side

will always equal the other side.